### v1

# 2INFN, Laboratori Nazionali del Gran Sasso, I-67100 Assergi, Italy 3Universität Heidelberg, Zentrum für Astronomie (ZAH), Institut für Theoretische Astrophysik, Albert Ueberle Str. 2, 69120, Heidelberg, Germany 4Universität Heidelberg, Interdisziplinäres Zentrum für Wissenschaftliches Rechnen, Heidelberg, Germany

11institutetext: 1Gran Sasso Science Institute (GSSI), Viale Francesco Crispi 7, 67100, L’Aquila, Italy
2INFN, Laboratori Nazionali del Gran Sasso, I-67100 Assergi, Italy
3Universität Heidelberg, Zentrum für Astronomie (ZAH), Institut für Theoretische Astrophysik, Albert Ueberle Str. 2, 69120, Heidelberg, Germany
4Universität Heidelberg, Interdisziplinäres Zentrum für Wissenschaftliches Rechnen, Heidelberg, Germany
5Physics and Astronomy Department Galileo Galilei, University of Padova, Vicolo dell’Osservatorio 3, I–35122, Padova, Italy
6INFN - Padova, Via Marzolo 8, I–35131 Padova, Italy
7INAF - Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, I-35122 Padova, Italy
8INAF Osservatorio Astronomico d’Abruzzo, Via Maggini, 64100 Teramo, Italy
9 Departament de Física Quàntica i Astrofísica, Institut de Ciències del Cosmos, Universitat de Barcelona, Martí i Franquès 1, E-08028 Barcelona, Spain
10LIGO Laboratory, Massachusetts Institute of Technology, Cambridge, MA 02139, USA
11Kavli Institute for Astrophysics and Space Research, Massachusetts Institute of Technology, Cambridge, MA 02139, USA
12Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, USA
13Elizabeth S. and Richard M. Cashin Fellow at the Radcliffe Institute for Advanced Studies at Harvard University, 10 Garden Street, Cambridge, MA 02138, USA
Black Hole - Neutron Star and Binary Neutron Star Mergers from Population III and II stars
Benedetta Mestichelli [Uncaptioned image] benedetta.mestichelli@gssi.it112233    Michela Mapelli [Uncaptioned image] mapelli@uni-heidelberg.de3344556677    Filippo Santoliquido[Uncaptioned image] 1122    Manuel Arca Sedda[Uncaptioned image] 112288
Marica Branchesi
1188
   Lavinia Paiella 11 8 8    Guglielmo Costa 5577    Giuliano Iorio 99    Matthew Mould [Uncaptioned image] 10101111    Veronika Lipatova [Uncaptioned image] 33    Boyuan Liu [Uncaptioned image] 33    Ralf S. Klessen [Uncaptioned image] 334412121313
Population III (Pop. III) stars are expected to be massive and to undergo minimal mass loss due to their lack of metals, making them ideal progenitors of black holes and neutron stars. Here, we investigate the formation and properties of binary neutron star (BNS) and black hole-neutron star (BHNS) mergers originating from Pop. III stars, and compare them to their metal-enriched Population II (Pop. II) counterparts, focusing on their merger rate densities (MRDs), primary masses and delay times. We find that, despite the high merger efficiency of Pop. III BNSs and BHNSs, their low star formation rate results in a MRD at least one order of magnitude lower than that of Pop. II stars. The MRD of Pop. III BNSs peaks at redshift $z\sim 15$, attaining a value $\mathcal{R}_{\rm BNS}(z\sim 15)\sim 15\,\rm Gpc^{-3}\,yr^{-1}$, while the MRD of Pop. III BHNSs is maximum at $z\sim 13$, reaching a value $\mathcal{R}_{\rm BHNS}(z\sim 13)\sim 2\,\rm Gpc^{-3}\,yr^{-1}$. Finally, we observe that the black hole masses of Pop. III BHNS mergers have a nearly flat distribution with a peak at $\sim 20\,{\rm M}_{\odot}$ and extending up to $\sim 50\,{\rm M}_{\odot}$. Black holes in Pop. II BHNS mergers show instead a peak at $\lesssim 15\,{\rm M}_{\odot}$. We consider these predictions in light of recent gravitational-wave observations in the local Universe, finding that a Pop. III origin is preferred relative to Pop. II for some events.
1 Introduction
Population III (Pop. III) stars formed from metal-free gas at high redshift (Haiman et al. 1996; Tegmark et al. 1997; Yoshida et al. 2003). No conclusive evidence of Pop. III stars exists so far (Rydberg et al. 2013; Schauer et al. 2022; Larkin et al. 2023; Meena et al. 2023; Trussler et al. 2023). Current models suggest that Pop. III stars have a top-heavier initial mass function (IMF) compared to stellar populations in the local Universe (Stacy & Bromm 2013; Susa et al. 2014; Hirano et al. 2015; Jaacks et al. 2019; Sharda et al. 2020; Liu & Bromm 2020a, b; Wollenberg et al. 2020; Chon et al. 2021; Tanikawa et al. 2021b; Jaura et al. 2022; Prole et al. 2022; Klessen & Glover 2023; Liu et al. 2024a). Additionally, because of their lack of metals, we expect these stars to lose a small amount of mass through stellar winds over their lifetimes, when they are not fast rotators (Heger et al. 2002; Kinugawa et al. 2014, 2016; Hartwig et al. 2016; Belczynski et al. 2017; Liu et al. 2021; Tanikawa et al. 2021a, 2022, 2023; Costa et al. 2023; Santoliquido et al. 2023; Tanikawa 2024).
In contrast, early Population II (Pop. II) stars are enriched by the yields of Pop. III stars, but still sufficiently metal poor (metallicity $Z$ up to a few ×$\times{}10^{-4}$, Choudhury et al. 2018; Frebel et al. 2019) that mass loss via line-driven stellar winds is quenched (Kudritzki & Puls 2000; Vink et al. 2001; Sabhahit et al. 2023).
Compact binary mergers from Pop. III star progenitors could be detected by current gravitational-wave (GW) interferometers, if their delay times are sufficiently long. As a matter of fact, one of the possible scenarios for the massive binary black hole (BH) merger GW190521, forming a remnant BH of $\sim 150\,{\rm M}_{\odot}$, is the collapse of a Pop. III binary. Indeed, the negligible mass loss and top-heavy IMF of Pop. III stars might lead to the formation of a population of massive BH and binary BHs (Abbott et al. 2019a, 2020a, 2020b; Liu & Bromm 2020b; Kinugawa et al. 2021; Tanikawa et al. 2021a). The properties of these massive compact objects have been extensively studied through a variety of numerical simulations both in isolation (Tanikawa et al. 2021b, a, 2022, 2023; Costa et al. 2023; Santoliquido et al. 2023; Tanikawa 2024) and in star clusters (Sakurai et al. 2017; Wang et al. 2022; Mestichelli et al. 2024; Liu et al. 2024b; Reinoso et al. 2025; Wu et al. 2025). Furthermore, binary BH mergers from Pop. III stars are expected to be significant sources for third-generation interferometers, such as Cosmic Explorer (Reitze et al. 2019; Evans et al. 2021, 2023), and the Einstein Telescope (Kalogera et al. 2021; Branchesi et al. 2023; Santoliquido et al. 2023, 2024; Abac et al. 2025), as these instruments will be capable of probing the high-redshift Universe, where Pop. III star formation is expected to be prevalent.
Besides binary BHs, Pop. III stars might also be the progenitors of binary neutron stars (BNSs) and BH-neutron star binaries (BHNSs) detectable with current and next-generation interferometers. By the end of the third observation run, the LIGO–Virgo–KAGRA (LVK) collaboration has observed several BHNS and BNS merger candidates (Abbott et al. 2019b, 2020c, 2021, 2023a, 2024). Furthermore, the merger GW230529, detected at the beginning of the fourth observing run, has primary component mass $\sim{3-6}$ M (Abac et al. 2024), inside the previously claimed mass gap between the maximum neutron star and minimum BH mass (Orosz 2003; Özel et al. 2010; Farr et al. 2011).
While current GW detectors only probe BNS and BHNS mergers in the nearby Universe (up to $z\sim 0.2$, Abbott et al. 2023a), next generation interferometers will be sensitive to neutron star mergers in a much larger portion of the Universe, increasing the likelihood of observing BNS and BHNS mergers from Pop. II and III star progenitors (Branchesi et al. 2023; Gupta et al. 2024; Abac et al. 2025).
Here, we present models of BNSs and BHNSs formed from the evolution of Pop. III and Pop. II binary stars. We study the distribution of their masses and delay times, and their merger rate densities. To encompass the uncertainties on Pop. III stars, we probe different combinations of IMFs and orbital parameters, following the same approach as our previous works on binary BHs (Costa et al. 2023; Santoliquido et al. 2023; Mestichelli et al. 2024). The paper is structured as follows. Section 2 presents the initial conditions and the numerical codes used. Section 3 shows our main results. Section 4 provides a more detailed analysis of the properties of BHNS mergers and compares formation scenarios involving Pop. III and Pop. II progenitors for currently observed LVK events using Bayes factors. Finally, Section 5 summarizes our conclusions.
2 Methods
2.1 Population synthesis with sevn
We model the formation of BNSs and BHNSs with the binary population synthesis code sevn(Spera et al. 2019; Mapelli et al. 2020; Iorio et al. 2023), which calculates single stellar evolution by interpolating a set of pre-computed single stellar tracks and describes the main binary interaction processes (mass transfer via stellar winds, Roche lobe overflow, common envelope, tides, and GW decay) by means of semi-analytic prescriptions (Hurley et al. 2002; Iorio et al. 2023). For the common-envelope phase, we set the efficiency parameter $\alpha_{\rm CE}=1$ and model the envelope binding energy $\lambda_{\rm CE}$ as done by Claeys et al. (2014). We refer to Costa et al. (2023), Santoliquido et al. (2023) and Mestichelli et al. (2024) for further details about our models.
2.2 Compact objects formation
According to the prescriptions in sevn , depending on the final carbon-oxygen (CO) core mass ($M_{\rm CO,f}$), a star can form a white dwarf ($M_{\rm CO,f}<1.38\,\rm M_{\odot}$), undergo electron-capture supernova and form a neutron star ($1.38\leq M_{\rm CO,f}<1.44\,\rm M_{\odot}$; Giacobbo & Mapelli 2019), or a core-collapse supernova and form either a neutron star or a BH (CCSN; $M_{\rm CO,f}\geq 1.44\,\rm M_{\odot}$). For CCSNe, we adopt the rapid formalism by Fryer et al. (2012), according to which if $6\leq M_{\rm CO,f}<7\,\rm M_{\odot}$ or $M_{\rm CO,f}\geq 11\,\rm M_{\odot}$, the star will collapse directly into a BH. As a consequence of the adopted CCSN model, neutron star masses fall between $1.19$ and $2\,{\rm M}_{\odot}$, and a lower mass gap between $2$ and $5\,{\rm M}_{\odot}$ naturally emerges. Given the lack of a predictive model for the mass function of neutron stars able to match the masses of Galactic neutron stars (Sgalletta et al. 2023), we do not include neutron star masses in our analysis, for both Pop. III and Pop. II progenitors.
For (pulsational) pair-instability supernovae, we used the fitting formulas reported by Mapelli et al. (2020) and based on the simulations by Woosley (2017). If the pre-supernova He-core mass of a star, $M_{\rm He,f}$, is between 32 and 64 ${\rm M}_{\odot}$, it undergoes pulsational pair-instability. A pair-instability supernova is triggered for $64\leq M_{\rm He,f}\leq 135\,{\rm M}_{\odot}$. Above $135\,{\rm M}_{\odot}$, the star collapses directly to a BH as a consequence of photodisintegration (Woosley 2017; Spera & Mapelli 2017; Renzo et al. 2020).
2.3 Supernova kicks
After a supernova (SN), compact objects receive a natal kick, that we model following Giacobbo & Mapelli (2020)
$v_{\rm kick}=f_{\rm H05}\,\frac{\langle M_{\rm NS}\rangle}{M_{\rm rem}}\,\frac{M_{\rm ej}}{\langle M_{\rm ej}\rangle},$ (1)
with $\langle M_{\rm NS}\rangle$ and $\langle M_{\rm ej}\rangle$ computed from a population of isolated neutron stars with metallicities representative of the Milky Way, $M_{\rm rem}$ the mass of the compact object, and $M_{\rm ej}$ the ejected mass. $f_{\rm H05}$ is a number randomly drawn from a Maxwellian distribution with one-dimensional root mean square $\sigma_{\rm kick}=265\,\mathrm{km\,s^{-1}}$, derived from the proper motions of young Galactic pulsars (Hobbs et al. 2005). This empirical formalism, while approximated, allows us to match the observations of young Galactic pulsars (Hobbs et al. 2005; Verbunt et al. 2017), and results in lower kicks for BHs, in accordance with observations (Atri et al. 2019). In particular, it implies that BHs forming through direct collapse receive a null natal kick.
In a binary system, natal kicks affect the orbital properties, relative orbital velocity, and center of mass of the binary (Hurley et al. 2002). After the kick, the orbital parameters are updated taking into account the new relative orbital velocity and total mass of the binary. For further details, we refer to Iorio et al. (2023).
2.4 Stellar tracks
Our Pop. III star tracks do not include rotation and are calculated with the stellar-evolution code parsec (Bressan et al. 2012; Chen et al. 2015; Costa et al. 2019, 2025). At the beginning of their main sequence, Pop. III stars cannot ignite the carbon-nitrogen-oxygen (CNO) tri-cycle because of the initial lack of these elements. In order to contrast the gravitational collapse with the energy provided by the proton-proton (pp) chain, Pop. III stars need to reach very high central temperatures. The temperature can become so high that some carbon is synthesized via triple – $\alpha$ reaction (He burning), even during the main sequence. As a consequence, the CNO tri-cycle ignites and replaces the pp chain as the main source of energy of the star (Marigo et al. 2001; Murphy et al. 2021). At the end of the main sequence, Pop. III stars have high enough central temperatures to transition smoothly to the core-He burning phase. Usually, these stellar evolution features appear at $Z<10^{-10}$ (Cassisi & Castellani 1993). As a consequence, for Pop. III stars we assume an initial hydrogen abundance $X=0.751$, He abundance $Y=0.2485$ (Komatsu et al. 2011), and metallicity $Z=10^{-11}$ (Tanikawa et al. 2021b; Costa et al. 2023). For Pop. II stars we adopt a metallicity $Z=10^{-4}$.
We consider stars with zero-age main sequence (ZAMS) mass $m_{\rm ZAMS}$ between 2 and 600 ${\rm M}_{\odot}$. Stars with $2\leq m_{\rm{ZAMS}}<10\,\rm M_{\odot}$ reach the early-asymptotic giant branch phase, while stars with $m_{\rm{ZAMS}}\geq 10\,\rm M_{\odot}$ evolve until late phases of core-oxygen burning, or the beginning of the pair-instability regime. All tracks are computed with the same setup as Costa et al. (2021), including stellar winds, nuclear reaction network, opacity, and equation of state. Above the convective core, we assume a penetrative overshooting with a characteristic parameter of $\lambda_{\rm ov}=0.5$ in units of pressure scale height. For further information on the evolutionary tracks, we refer to Costa et al. (2023).
In Figure 1 we show the evolution of the radii of Pop. III and Pop. II stars. This Figure highlights the compactness of metal-free stars compared to Pop. II stars, especially in the mass range $8\leq m_{\rm ZAMS}\leq 30\,{\rm M}_{\odot}$. As already pointed out by Costa et al. (2023), Pop. III stars with $m_{\rm ZAMS}<100\,{\rm M}_{\odot}$ begin burning He in their cores before their Pop. II counterparts. This difference becomes less evident at higher masses. We also notice that Pop. III stars with $m_{\rm ZAMS}>100\,{\rm M}_{\odot}$ reach the pre-SN stage as blue supergiant stars, while Pop. II stars explode as very large red supergiant stars.
Figure 1: Evolution of stellar radii for $8\leq m_{\rm ZAMS}\leq 300\,{\rm M}_{\odot}$. We represent Pop. III stars with a solid black line and Pop. II stars with a dashed red line. The orange stars (circles) represent the beginning (end) of core-He burning. The blue circles indicate instead the pre-SN phase.
2.5 Binary initial conditions
The simulation sets we analyze here are the same as those presented by Costa et al. (2023). In this work, we analyze BNSs and BHNSs for the first time, whereas Costa et al. (2023) focused on binary BHs.
We assume the following fiducial models for the IMFs and initial orbital properties of the simulated binary stars. For both Pop. II and Pop. III binary stars, we randomly sample the initial mass ratios, orbital periods and eccentricities from the distributions reported by Sana et al. (2012) and obtained from observations of Galactic O-type binary stars. For Pop. III binary stars, we randomly sample the ZAMS masses of Pop. III primaries from a log-flat IMF, as motivated by previous studies (Stacy & Bromm 2013; Susa et al. 2014; Hirano et al. 2015; Wollenberg et al. 2020; Chon et al. 2021; Tanikawa et al. 2021a; Jaura et al. 2022; Prole et al. 2022). In contrast, we randomly sample the ZAMS masses of Pop. II primaries from a Kroupa IMF (Kroupa 2001). Hereafter, we refer to our fiducial model for Pop. III and Pop. II binary stars as log1 and kro1, respectively, for consistency with previous works (Costa et al. 2023; Santoliquido et al. 2023; Mestichelli et al. 2024).
In Sections 3.1.1, 3.2.1, and 4.1, we compare our fiducial models with additional simulations that assume a variety of initial conditions, encompassing the uncertainties on both IMF and orbital properties of Pop. III binary stars. We describe the additional models in Appendix A and summarize them in Table 1. We wish to point out that in this work we do not explore the possibility of a minimum ZAMS mass smaller than $2\,{\rm M}_{\odot}$, and, as a consequence, we do not analyze the impact of stars with $m_{\rm ZAMS}<2\,{\rm M}_{\odot}$.
Model $m_{\mathrm{ZAMS,1}}$ $m_{\mathrm{ZAMS}}$ $q$ $P$ $e$
kro1 Kroupa (2001) Flat in log S12 S12 S12
kro5 Kroupa (2001) Flat in log SB13 SB13 Thermal
lar1 Larson (1998) - S12 S12 S12
lar5 Larson (1998) - SB13 SB13 Thermal
log1 Flat in log - S12 S12 S12
log2 Flat in log - S12 SB13 Thermal
log3 - - Sorted S12 S12
log4 - - SB13 S12 Thermal
log5 Flat in log - SB13 SB13 Thermal
top1 Top heavy - S12 S12 S12
top5 Top heavy Flat in log SB13 SB13 Thermal
Table 1: Initial conditions for the stellar population models analyzed in this work. Column 1: name of the model. Column 2: IMF for the primary star. Column 3: ZAMS mass of the overall stellar population. Columns 4, 5 and 6: initial distributions for the mass ratio $q$, period $P$ and eccentricity $e$. The acronyms S12 and SB13 refer to Sana et al. (2012) and Stacy & Bromm (2013), respectively.
2.6 Formation channels
To interpret the properties of the simulated BNS and BHNS mergers, we will refer to the six formation channels introduced by Iorio et al. (2023) and defined as follows. Channel 1 involves stable mass transfer before the formation of the first compact object, followed by at least one common-envelope phase. In channel 2, the binary evolves only through stable mass transfer episodes. In both channels 3 and 4, the binary undergoes at least one common-envelope phase before the formation of the first compact object. In particular, in channel 3, the companion star retains an H-rich envelope at the time of formation of the first compact object. In channel 4, instead, the companion has been stripped of its envelope when the first compact remnant forms. The two remaining channels refer to systems that do not interact at all (channel 0), or interact only after the formation of the first compact object (channel 5). A summary of these formation channels can be found in Table 2.
Channel Description
0 No interaction
1 SMT before first remnant + CE after first remnant
2 SMT only
3 CE before first remnant + secondary with H-rich envelope at formation of first remnant
4 CE before first remnant + secondary stripped of envelope at formation of first remnant
5 Interaction only after formation of first remnant
Table 2: Formation channels of merging compact binaries. Column 1: number of the channel. Column 2: brief description of the formation channel. Here SMT stands for stable mass transfer, and CE stands for common envelope.
2.7 Merger rate density
We estimate the merger rate density (MRD) of compact binaries using the semi-analytic code $\mathcal{R}$ate (Santoliquido et al. 2020, 2021, 2023), which combines catalogs of simulated compact binary mergers with a model for the metal-dependent cosmic star formation rate density. Specifically, the MRD in the comoving frame is computed as:
$\mathcal{R}(z)=\int_{z_{{\rm{max}}}}^{z}\left[\int_{Z_{{\rm{min}}}}^{Z_{{\rm{max}}}}\,{}\psi{(z^{\prime},Z)}\,{}\mathcal{F}(z^{\prime},z,Z)\,{}{\rm{d}}Z\right]\,{}\frac{{{\rm d}t(z^{\prime})}}{{\rm{d}}z^{\prime}}\,{}{\rm{d}}z^{\prime},$ (2)
where $\psi(z^{\prime},Z)$ is the metallicity-dependent cosmic star formation rate density at redshift $z^{\prime}$ and metallicity $Z$, and ${{\rm{d}}t(z^{\prime})}/{{\rm{d}}z^{\prime}}=H_{0}^{-1}\,{}(1+z^{\prime})^{-1}\,{}[(1+z^{\prime})^{3}\Omega_{M}+\Omega_{\Lambda}{}]^{-1/2}$, where $H_{0}$ is the Hubble parameter and $\Omega_{M}$ and $\Omega_{\Lambda}$ are the matter and energy density, respectively. We adopt here the values of the cosmological parameters from Planck Collaboration et al. (2020).
The term $\mathcal{F}(z^{\prime},z,Z)$ in Equation 2 is given by:
$\mathcal{F}(z^{\prime},z,Z)=\frac{1}{M_{\ast{}}(z^{\prime},Z)}\,{}\frac{{\rm{d}}\mathcal{N}(z^{\prime},z,Z)}{{\rm{d}}t(z)},$ (3)
where ${{\rm{d}}\mathcal{N}(z^{\prime},z,Z)/{\rm{d}}}t(z)$ is the rate of BNS or BHNS mergers from stars with metallicity $Z$ that form at redshift $z^{\prime}$ and merge at redshift $z$, and $M_{\ast}(z^{\prime},Z)$ is the initial total stellar mass that forms at redshift $z^{\prime}$ with metallicity $Z$. We obtain ${{\rm{d}}\mathcal{N}(z^{\prime},z,Z)/{\rm{d}}}t(z)$ directly from our simulations.
Owing to the lack of direct observations of Pop. III stars, numerous models have been proposed for their $\psi(z^{\prime},Z)$. We adopt the star formation history derived from the semi-analytic model a-sloth, developed by Hartwig et al. (2022) and Magg et al. (2022), which tracks individual Pop. III and Pop. II stars and is calibrated against a range of observables from both the local and high-redshift Universe. For comparison, we also estimate the MRD of BNSs and BHNSs formed from early Pop. II stars. In this case, we adopt the star formation rate density of Pop. II stars ($Z\sim 10^{-5}-10^{-2}\,\rm Z_{\odot}$) by Liu et al., in prep., obtained adopting a revised version of the semi-analytic code a-sloth (Hartwig et al. 2024). This version is calibrated to reproduce the metallicity – stellar mass – star formation rate relation of high-redshift galaxies in recent observations of the James Webb Space Telescope (Sarkar et al. 2025).
The described star formation rate densities $\psi(z^{\prime},Z)$ are reported in Fig. 2. The adopted models for $\psi(z^{\prime},Z)$ predicted by a-sloth are valid for $z>4.5$. Therefore, it is assumed that Pop. III and Pop. II star formation at later epochs is negligible. This is supported by the finding that their star formation rate density peaks at $z\sim 10-15$, and has dropped by a factor of 5 - 50 by $z\sim 4.5$, because of metal enrichment and re-ionization. We also show the uncertainty range associated with Pop. II star formation, reflecting variations in SN-driven outflows and assumptions on the IMF (Liu et al., in prep.)
Figure 2: Star formation rate density of Pop. III (magenta dash-dotted line, Hartwig et al. 2022) and Pop. II (purple dashed line, Liu et al., in prep.) stars. The shaded purple region denotes the uncertainty range associated with Pop. II star formation history. As a comparison, the black solid line shows the star formation rate density of Pop. II-I stars by Madau & Fragos (2017).
3 Results
3.1 Black hole – neutron star binaries (BHNSs)
In this Section, we describe the properties of BHNS mergers from Pop. III and Pop. II binary stars. We first discuss the impact of the simulated Pop. III models on the MRD of BHNSs and make a comparison with the MRD of Pop. II BHNSs. We then explore the distributions of delay times, primary masses and mass ratios of BHNSs for the fiducial models of Pop. III and Pop. II star progenitors.
3.1.1 Merger rate density
Figure 3 shows the MRD of BHNSs. Here, we consider all the models summarized in Table 1 for Pop. III stars. For Pop. II stars, instead, we take into account only the fiducial model (kro1). As a comparison, we also show the total MRD of Pop. II-I BHNSs reported by Iorio et al. (2023) (see Appendix B for details).
The MRD of BHNSs from Pop. III stars peaks at $z\sim 13$, where the assumed models span about three orders of magnitude $\left[\mathcal{R}_{\rm BHNS}(z\sim 13)\sim 10^{-2}-2\,\rm Gpc^{-3}\,yr^{-1}\right]$. The position of the peak is the result of a convolution between the chosen Pop. III star formation rate density (Hartwig et al. 2022) and the predominantly short delay times of BHNS mergers (see Sec. 3.1.2 and Fig. 4). At the peak, the largest MRD is associated with model lar1 (see Table 1 and Appendix A). Model lar1 produces a $\mathcal{R}_{\rm BHNS}(z\sim 13)\sim 1.7\,\rm Gpc^{-3}\,yr^{-1}$, approximately four times larger than the MRD peak of our fiducial model ($\mathcal{R}_{\rm BHNS}(z\sim 13)\sim 0.4\,\rm Gpc^{-3}\,yr^{-1}$). In fact, although top-heavy IMFs tend to produce more BH and neutron star progenitors, the expansion of these massive progenitors during their evolution very often leads to premature mergers of the stellar components, inhibiting the formation of a compact binary (see Fig. 1). At $z=0$, all the Pop. III star models produce MRDs between $\sim 10^{-3}$ and $\sim 10^{-1}\,\rm Gpc^{-3}\,yr^{-1}$. In agreement with Santoliquido et al. (2023), we find that models with initial orbital periods from Sana et al. (2012) produce larger MRDs than models relying on Stacy & Bromm (2013) (log2, log5, kro5, lar5, top5). In fact, the former have smaller orbital periods than the latter, favoring BHNS mergers.
The MRD of Pop. II BHNSs peaks at $z\sim 9$ reaching $\mathcal{R}_{\rm BHNS}(z\sim 9)\sim 14\,\rm Gpc^{-3}\,yr^{-1}$. At $z=0$, Pop. II BHNSs yield a MRD approximately seven times higher than that of Pop. III stars under model lar1. The MRD from Pop. II BHNSs continues to dominate over that of Pop. III stars up to $z\sim 17$, even under the most favorable assumptions for Pop. III.
Finally, Pop. II-I BHNSs yield a MRD peaking at $z\sim 2.5$ ($\mathcal{R}_{\rm BHNS}(z\sim 2.5)\sim 60\,\rm Gpc^{-3}\,yr^{-1}$) due to the combined effect of star formation rate density (Madau & Fragos 2017) and dependence on metallicity (Iorio et al. 2023). Our comparison sample from Iorio et al. (2023) including Pop. I stars dominates over the contribution of Pop. II stars up to $z\sim 7$.
Here, we note the seeming contradiction that the BHNS MRD from Pop. II-I stars is lower than the BHNS MRD from Pop. II stars alone at $z\sim{7-15}$. However, we remind that the compact object MRD from Pop. II-I stars reported by Iorio et al. (2023) was obtained with several different assumptions compared to the new MRDs presented here. The main differences are the maximum ZAMS mass (assumed to be only 150 M by Iorio et al. 2023), and the different cosmic star formation rate density $\psi{}(z,Z)$. Specifically, Iorio et al. (2023) obtain $\psi{}(z,Z)$ from Madau & Fragos (2017) assuming a log-normal metallicity distribution with a narrow dispersion $\sigma_{\mathrm{Z}}=0.1$ (see Fig. 2 and Appendix B), whereas here we assume self-consistent values of $\psi{}(z,Z)$ for both Pop. II and III stars directly derived from a-sloth (Hartwig et al. 2022, 2024; Liu at al., in prep.). Assuming a narrow metallicity dispersion suppresses pockets of metal-poor star formation, as already discussed by Sgalletta et al. (2024).
Table 3 reports a summary of the main results of this Section.
Model Population $\mathcal{R}_{\rm BHNS}(z=0)$ $\mathcal{R}_{\rm BHNS}(z=13.2)$ $\mathcal{R}_{\rm BHNS}(z=8.9)$
[$\rm Gpc^{-3}\,yr^{-1}$] [$\rm Gpc^{-3}\,yr^{-1}$] [$\rm Gpc^{-3}\,yr^{-1}$]
lar1 Pop. III $3.5\times 10^{-2}$ 1.7 0.8
kro1 Pop. II 0.2 6.5 13.7
Table 3: Merger rate density of BHNSs from Pop. III and Pop. II stars. Column 1: name of the model producing the largest merger rate density. In the case of Pop. II stars, only the fiducial model kro1 was considered. Column 2: stellar population. Column 3: MRD at $z=0$. Column 4: MRD at the peak of the Pop. III distribution. Column 5: MRD at the peak of the Pop. II distribution.
Figure 3: Merger rate density of BHNSs as a function of the redshift $z$. The colored lines show the distributions for all the simulated Pop. III models. The thick solid blue line represents our fiducial model for Pop. III stars (log1). The thick solid gray line is the fiducial model for Pop. II stars (kro1); the gray shaded area encompasses the uncertainties on galactic outflows and IMF parameters (see Sec. 2.7). Finally, the thick solid black line represents the merger rate density of Pop. II-I BHNSs estimated by Iorio et al. (2023).
Model Population $\eta_{\rm BHNS}$ Ch. 0 Ch. 1 Ch. 2 Ch. 3 Ch. 4 Ch. 5
[${\rm M}_{\odot}^{-1}$] (%) (%) (%) (%) (%) (%)
log1 Pop. III $1.2\times 10^{-5}$ 0.6 7.0 55.3 29.0 0 8.0
kro1 Pop. II $3.4\times 10^{-5}$ 0 44.8 46.9 1.4 0.2 6.7
Table 4: Formation channels of BHNS mergers. Column 1: name of the model. Column 2: stellar population. Column 3: merger efficiency of BHNSs $\eta_{\rm BHNS}=N_{\rm BHNS,m}/M_{*}$, with $N_{\rm BHNS,m}$ the number of BHNS mergers. Column 4-9: percentage of BHNS mergers happening through channel 0-5.
3.1.2 Delay times and formation channels
Figure 4 shows the delay time (i.e., the time between the formation of the stellar binary system and its merger) for BHNS mergers in our fiducial Pop. III and Pop. II models. We compare these distributions with the $\propto t^{-1}$ trend from Dominik et al. (2012). This is the expected scaling of $t_{\rm del}$ if we consider orbital periods with a nearly log-uniform distribution (as in kro1 and log1, see Table 1). Deviations from this trend suggest a final semi-major axis distribution significantly different compared to the initial one (Peters 1964; Sana et al. 2012). In general, we find a good agreement between our simulations and the $\propto t^{-1}$ trend.
We point out a deviation from the trend for Pop. III BHNS mergers with $t_{\rm del}>10^{3}\,\rm Myr$, which is due to a majority of systems with semi-major axes larger than $\sim 10\,\rm R_{\odot}$. In Table 4, we report the fraction of BHNS mergers happening through the formation channels previously described in Sec. 2.6, and summarized in Table 2. We find that Pop. III BHNSs predominantly form via channel 2, which involves only stable mass transfer episodes and, as a result, does not significantly reduce the semi-major axis.
By contrast, Pop. II BHNS mergers show an excess at short delay times ($t_{\rm del}\sim 15\,\rm Myr$). This is due to the dominance of Pop. II BHNSs forming via channel 1, which shrinks semi-major axes effectively via common envelope.
Figure 4: Distribution of delay times of BHNS mergers for the fiducial models of Pop. III (continuous blue line) and Pop. II stars (dotted blue line). The gray dashed line shows the $\propto t^{-1}$ trend from Dominik et al. (2012).
3.1.3 Primary mass and mass ratio
Figure 5 shows the primary mass (i.e., the BH mass) distribution of BHNS mergers. Pop. II BHNS mergers show a peak at $m_{1}\lesssim 15\,{\rm M}_{\odot}$ (Giacobbo & Mapelli 2018; Broekgaarden et al. 2021). After this peak, the number of BHNS mergers decreases with $m_{1}$ up to $\sim 60\,{\rm M}_{\odot}$. BHNS mergers from Pop. III stars, instead, have a flatter distribution in $m_{1}$ between 5 and $\sim 47\,{\rm M}_{\odot}$ and show a peak around $22\,{\rm M}_{\odot}$. The different trends of these distributions are related to the dependence of the mass-radius relation on the metallicity $Z$ of the stars. In fact, regardless of the formation channel of BHNS mergers, Pop. II star binaries with $m_{1,\rm ZAMS}>40\,{\rm M}_{\odot}$ will tend to prematurely merge unless $a_{\rm ZAMS}\gtrsim 10^{3}\,\rm R_{\odot}$. Pop. III star binaries with high $m_{1,\rm ZAMS}$ will instead expand less during their evolution, and avoid premature mergers if $a_{\rm ZAMS}\gtrsim 400\,\rm R_{\odot}$. As reported in Table 4, the lack of a prominent peak at $m_{1}<20\,{\rm M}_{\odot}$ and the choice of a top-heavier IMF, lead to a lower merger efficiency for Pop. III BHNSs $\eta_{\rm BHNS}$.
Figure 6 shows the dependence of the primary mass of BHNS mergers on the delay time $t_{\rm del}$. In the case of Pop. III stars, high-mass systems merge preferentially at large delay times. This means that, at low-redshift, we might be able to observe GW signals from BHNSs with $m_{1}>20\,{\rm M}_{\odot}$ deriving from the evolution of metal-free binaries. In fact, more than $85\%$ of Pop. III BHNS mergers have $m_{1}>20\,{\rm M}_{\odot}$, while the fraction reduces to $\sim 11\%$ for Pop. II. As a consequence, even though at low redshifts the MRD of Pop. II BHNSs is about one order of magnitude larger than the one of Pop. III BHNSs (see Fig. 3), we have eight times more mergers from Pop. III BHNSs with $m_{1}>20\,{\rm M}_{\odot}$ compared to Pop. II. The contribution of Pop. III BHNS mergers to the currently observed GW events will be further discussed in Sec. 4.2.
Similarly to Pop. III, Pop. II BHNSs with massive primary BHs tend to merge with longer delay times (bottom panel of Figure 6). In fact, Pop. II binaries with large $m_{\rm ZAMS,1}$ are generally able to avoid premature mergers only when their initial semi-major axis is wide enough, resulting in longer delay times.
Figure 5: Distribution of primary mass of BHNS mergers for the fiducial models of Pop. III (continuous blue line) and Pop. II (dotted blue line).
Figure 6: Distribution of primary mass of BHNS mergers in three ranges of $t_{\rm del}$. We show the results for the fiducial models of Pop. III (top), and Pop. II progenitors (bottom). We represent in violet mergers with $t_{\rm del}<20\,\rm Myr$, in blue-green mergers with $20<t_{\rm del}<10^{3}\,\rm Myr$, and in green mergers with $t_{\rm del}>10^{3}\,\rm Myr$.
Finally, Fig. 7 shows the distributions of mass ratio $q$ of the simulated BHNS mergers. Pop. III stars yield a distribution that peaks at $q<0.1$ and extends up to $q\sim 0.3$; in fact, as we have seen, metal-free stars favor BHNS mergers with a large primary mass, resulting in small values of $q$. In contrast, Pop. II BHNSs merge with a mass ratio that peaks at $q\sim 0.2$ and extends to larger values, up to $q\sim 0.35$. In accordance with both Broekgaarden et al. (2021) and Giacobbo & Mapelli (2018), this is a consequence of the peak at $m_{1}<15\,{\rm M}_{\odot}$ (Fig. 5) and of the lower mass gap, between $2$ and $5\,{\rm M}_{\odot}$. These distinguishing features are particularly interesting because they can as well contribute to a distinction between GW events from Pop. III and Pop. II binaries (see Sec. 4.2).
Figure 7: Distribution of mass ratios $q$ of BHNS mergers for the fiducial models of Pop. III (continuous blue line) and Pop. II (dotted blue line).
3.2 Binary neutron stars
In this Section, we present the properties of BNS mergers from Pop. III and Pop. II stars. Firstly, we display the MRD of BNSs from different simulated models of Pop. III stars and compare them to the MRD of Pop. II BNSs. We then discuss the delay time distributions.
3.2.1 Merger rate density
Figure 8 shows the MRD of BNSs from Pop. III and Pop. II stars. For Pop. III stars, we take into account all the simulated models reported in Table 1. In the case of Pop. II stars, instead, we only consider the fiducial model kro1. As we did for BHNSs, we compare these MRDs with the one derived by Iorio et al. (2023), which includes higher metallicity stars.
The MRD of BNSs from Pop. III stars peaks at $z\sim 15$, for all considered models. Here, the different models we simulated span five orders of magnitude $\left[\mathcal{R}_{\rm BNS}(z\sim 15)\sim 10^{-3}-15\,\rm Gpc^{-3}\,yr^{-1}\right]$. Because of the extremely short BNS delay times (see Sec. 3.2.2 and Fig. 9), the MRDs from Pop. III stars peak in the vicinity of the peak of the chosen star formation rate density. The trend of all Pop. III MRDs at $z<5$ is also due to the convolution of the small delay times ($<50\,\rm Myr$) with the chosen star formation rate density $\psi(z^{\prime},Z)$ (see Fig. 2). Model kro1 produces the highest MRD for Pop. III stars, with $\mathcal{R}_{\rm BNS}(z)\sim 15\,\rm Gpc^{-3}\,yr^{-1}$ at $z\sim 15$. The peak of our fiducial Pop. III model (log1) is an order of magnitude lower. In fact, kro1 produces less massive BNS progenitors involved in premature mergers. As discussed in Sec. 3.1.1, models adopting the initial orbital period distribution from Stacy & Bromm (2013) also result in lower MRDs in this case.
The BNS MRD from Pop. II stars dominates over the one from Pop. III stars up to $z\sim 10$, where it reaches a value of $\mathcal{R}_{\rm BNS}(z\sim 10)\sim 5\,\rm Gpc^{-3}\,yr^{-1}$. There, the MRD of Pop. III BNSs from model kro1 is within the same order of magnitude. As pointed out for Pop. III BNSs, also Pop. II BNS mergers have extremely short delay times, leading to a drop in the MRD at $z<5$.
Figure 1 shows that Pop. III stars with $m_{\rm ZAMS}\lesssim 20\,{\rm M}_{\odot}$ are more compact with respect to Pop. II stars. As a consequence, they produce, on the one hand, less premature mergers and, on the other, a higher BNS merger efficiency ($\eta_{\rm BNS}\sim 10^{-4}\,{\rm M}_{\odot}^{-1}$ for model kro1). However, as shown in Table 5, convolving the merger efficiency with the star formation rate density models yields a MRD from Pop. II BNSs that exceeds that of Pop. III by at least one order of magnitude at $z<10$.
For comparison, we show that the MRD of BNSs born from Pop. II-I progenitors (Iorio et al. 2023) peaks at $z\sim 2$ ($\mathcal{R}_{\rm BNS}(z\sim 2)\sim 10^{3}\,\rm Gpc^{-3}\,yr^{-1}$), consistently with the peak of the star formation rate density from Madau & Fragos (2017). The contribution of BNSs born from more metal-rich progenitors dominates the MRDs below $z\sim 11$.
Figure 8: Same as Fig. 3 but for BNS mergers.
Model Population $\mathcal{R}_{\rm BNS}(z=0)$ $\mathcal{R}_{\rm BNS}(z=15.2)$ $\mathcal{R}_{\rm BNS}(z=9.8)$
[$\rm Gpc^{-3}\,yr^{-1}$] [$\rm Gpc^{-3}\,yr^{-1}$] [$\rm Gpc^{-3}\,yr^{-1}$]
kro1 Pop. III $1.4\times 10^{-4}$ 15.4 3.2
kro1 Pop. II $2.2\times 10^{-3}$ 1.6 4.9
Table 5: Same as Table 3 but for the merger rate density of BNSs from Pop. III and Pop. II stars.
Model Population $\eta_{\rm BNS}$ Ch. 0 Ch. 1 Ch. 2 Ch. 3 Ch. 4 Ch. 5
[${\rm M}_{\odot}^{-1}$] (%) (%) (%) (%) (%) (%)
log1 Pop. III $6.2\times 10^{-6}$ 0 87.1 0 3.6 0.2 9.1
kro1 Pop. II $4.3\times 10^{-6}$ 0 45.4 0 53.5 0.4 0.7
Table 6: Same as Table 4 but for BNS mergers.
3.2.2 Delay times and formation channels
Figure 9 shows the distribution of delay times of BNS mergers for the fiducial models of Pop. III and Pop. II stars. Both populations result in distributions deviating significantly from the $\propto t^{-1}$ trend predicted by Dominik et al. (2012). The delay times of BNS mergers show a peak at $\sim 20\,\rm Myr$, and a common trend up to $t_{\rm del}\sim 10^{3}\,\rm Myr$, after which we observe an excess of BNS mergers from Pop. II stars.
The peak at $\sim 20\,\rm Myr$ is the result of very efficient orbital shrinking after one (or more) common envelope event(s). As a matter of fact, Table 6 shows that BNS mergers take place when a binary system undergoes at least one common envelope during its evolution (channels 1,3,4,5). As a consequence, more than $98\,\%$ of the merging BNSs have $a<1\,\rm R_{\odot}$.
The excess of Pop. II systems at $t_{\rm del}>10^{3}\,\rm Myr$ can be also explained by looking at Table 6. In fact, half of the BNS mergers from Pop. II stars form through channel 3 (see Table 2), which favors binaries with $a_{\rm ZAMS}\sim 200-5\times 10^{4}\,\rm R_{\odot}$. Indeed, Fig. 1 shows that Pop. II stars with $m_{\rm ZAMS}\sim 8-20\,{\rm M}_{\odot}$ tend to expand to larger radii with respect to Pop. III stars in the same mass range. Consequently, only binary stars with large initial separations will avoid to merge prematurely. We find that a fraction of the binaries evolving through channel 3 yields BNSs with $a>2\,\rm R_{\odot}$, that will need $\geq 10^{3}\,\rm Myr$ to merge.
Lastly, Fig. 9 points out the lower merger efficiency of Pop. II BNSs, with an evident lack of systems merging with $50<t_{\rm del}<10^{3}\,\rm Myr$. In Table 6 we show that, for the chosen fiducial models, the merger efficiency of Pop. II BNSs is $\sim 1.4$ times lower than the one of Pop. III BNSs.
Figure 9: Same as Fig. 4 but for BNS mergers.
4 Discussion
In the following, we discuss the impact of the initial conditions on the properties of BHNS mergers. We then compute the detectability-conditioned Bayes factors for three GW events comparing a Pop. III to a Pop. II origin.
4.1 Impact of initial conditions on distributions of $m_{1}$ and $q$
Figure 10 shows the distribution of the primary mass $m_{1}$ of BHNS mergers for all the simulated models of Pop. III stars (Table 1). As a comparison, we also show the distributions for the two Pop. II star models assuming a Kroupa IMF (kro1, kro5; Kroupa 2001).
All Pop. III models produce BHNS mergers with a primary mass between $\sim 6\,{\rm M}_{\odot}$ and $\sim 50\,{\rm M}_{\odot}$, with the exception of log3, which pushes $m_{1}$ above the upper mass gap, up to $\sim 400\,{\rm M}_{\odot}$. In fact, the initial mass-ratio distribution adopted in model log3 (”sorted”; see Table 1) favors the formation of binaries with $m_{1,\rm ZAMS}\gtrsim 200\,{\rm M}_{\odot}$ and $m_{2,\rm ZAMS}\lesssim 25\,{\rm M}_{\odot}$. These high-mass binaries typically have initial separations $a_{\rm ZAMS}>10^{3}\,\rm R_{\odot}$ and merge predominantly through channel 3 (Table 2), which favors unequal-mass systems (Costa et al. 2023).
Models with initial orbital parameters from Stacy & Bromm (2013) (log2, log5, kro5, lar5, top5) tend to produce BHNS mergers with $m_{1}\gtrsim 15\,{\rm M}_{\odot}$ and peaking at $m_{1}\gtrsim 20\,{\rm M}_{\odot}$. These models yield less BHNS mergers because of the wide initial semi-major axes; merging BHNSs have $a_{\rm ZAMS}\sim 10^{3}-10^{5}\,\rm R_{\odot}$ and evolve mainly through channel 3. Differently from log3, these models rely on an initial mass-ratio distribution from Stacy & Bromm (2013) with $q_{\rm min}=0.1$, which favors the formation of BHNSs with large $m_{\rm 1,ZAMS}$, but does not allow systems with $m_{\rm 1,ZAMS}>200\,{\rm M}_{\odot}$.
Finally, models based on Sana et al. (2012) yield a flatter $m_{1}$ distribution, with a sharp peak around $m_{1}\sim 22\,{\rm M}_{\odot}$. We point out that model kro1 produces a factor three more mergers with respect to the Pop. III fiducial model log1.
Consequently, all the simulated Pop. III star models show a nearly flat distribution of BH mass for $m_{1}>20\,{\rm M}_{\odot}$. Moreover, in Sec. 3.1.3, we have seen that these massive systems usually merge with large delay times. Thus, it is important to take into account the contribution of Pop. III massive BHNS mergers also at low redshifts (see Sec. 4.2).
Pop. II BHNS mergers from model kro1 show a peak at $m_{1}\sim 10\,{\rm M}_{\odot}$ and follow a decreasing trend up to $\sim 60\,{\rm M}_{\odot}$ (see Sec. 3.1.3).
As in the case of Pop. III, Pop. II BHNS mergers from model kro5 (Stacy & Bromm 2013) mostly have $m_{1}>15\,{\rm M}_{\odot}$. Model kro5 is thus able to reproduce massive BHNS mergers without the help of stellar dynamics (Rastello et al. 2020; Arca Sedda 2020; Arca sedda et al. 2024). Nevertheless, because of their larger radial expansion with respect with Pop. III stars, Pop. II binaries are also able to produce more BHNS mergers with low-mass primaries.
Figure 11 shows the mass-ratio $q$ distributions for all the simulated models. Starting from Pop. III stars, we notice that models with initial orbital periods from Sana et al. (2012) peak at $q\sim 0.1$ and extend up to $q\sim 0.3$. Models based on Stacy & Bromm (2013), instead, get to a maximum of $q\sim 0.2$. We point out again model log3, producing BHNS mergers with extremely low mass ratios; this feature was already found and discussed by Costa et al. (2023) in the case of binary BHs. Moving on to Pop. II BHNS mergers, we see that model kro1 produces a peak at $q\sim 0.2$ and extends to larger mass ratios with respect to its metal-free counterpart (see also Sec. 3.1.3). Due to its primary mass distribution, model kro5 shows intermediate features between Pop. III BHNS mergers with initial parameter distributions from Stacy & Bromm (2013) and from Sana et al. (2012).
Figure 10: Primary BH mass distribution of BHNS mergers. The left-hand panel shows the distributions for all the simulated models of Pop. III stars reported in Table 1. The right-hand panel shows the $m_{1}$ distributions of Pop. II stars for models kro1 and kro5.
Figure 11: Same as Fig. 10 but for the mass ratio $q$.
4.2 Model selection with GW events
Here, we compare our models by calculating the detectability-conditioned posterior odds for GW events with BHNS components. To carry out this comparison, we follow the formalism presented in Mould et al. (2023), which we briefly summarize in Appendix C. Selection effects are computed using public LVK injections (Abbott et al. 2023) with a detection threshold in the false-alarm rate $<0.25\,\mathrm{yr}^{-1}$ (Abbott et al. 2023b).
Table 7 presents the detectability-conditioned Bayes factors, $\ln\mathcal{D}_{A/B}$, where model $A$ corresponds to a Pop. III model (as listed in Table 1) and model $B$ represents the fiducial Pop. II model (kro1). These values are obtained by convolving the merger catalogs generated with sevn with $\mathcal{R}$ate, following the methodology described in Sec. 2.7, and considering only mergers occurring at redshifts $z<2$. We compute Bayes factors for three GW events: GW190917, GW191219, and GW200105 (Abbott et al. 2021, 2024, 2023a). Our analysis includes GW190917, with a secondary mass within the lower mass gap ($2-5\,{\rm M}_{\odot}$). The waveform parameters used to evaluate the Bayes factors are the primary mass, mass ratio, and redshift of each event. In this analysis, we only consider events with sufficient overlap between the observed parameters and the simulated distributions to ensure meaningful Bayes factor estimates. A decisive preference between models is typically indicated when $|\ln\mathcal{D}_{A/B}|\gtrsim 5$ (Jeffreys 1939).
Among the selected GW events, only one (GW191219) has primary mass $m_{1}>20\,{\rm M}_{\odot}$. It is generally expected that Pop. II BHNS mergers with such high primary masses form in star clusters (Rastello et al. 2020; Arca Sedda 2020; Arca sedda et al. 2024). However, as illustrated in Figs. 5 and 6, an alternative formation channel could derive from the evolution of isolated Pop. III binaries.
GW191219 favors a Pop. III origin. This system is characterized by a highly asymmetric mass ratio ($q\sim 0.04$), leading to higher Bayes factors for models that favor mergers with extremely low mass ratios, such as kro5, lar5, log3, log4, log5, and top5. This outcome is consistent with Fig. 11, which shows that all simulated Pop. III models predominantly yield BHNS mergers with $q<0.1$. It should be noted that the false alarm rate associated with GW191219 is $\sim 4\,\rm{yr}^{-1}$, which is not consistent with the threshold typically adopted for estimating selection effects in the calculation of detection-conditioned Bayes factors. Nevertheless, the detectability-free Bayes factors $\mathcal{B}_{\rm Pop.\leavevmode\nobreak\ III/Pop.\leavevmode\nobreak\ II}$ for this event are found to be in close agreement with the detectability-weighted values reported in the table.
The remaining two GW events, GW190917 and GW200105, both have primary masses $m_{1}<20\,{\rm M}_{\odot}$ and show stronger support for a Pop. II formation channel.
As a consequence, we find that Pop. III BHNS mergers contribute to the population of systems with high primary mass and low mass ratio at low redshift more than Pop. II. In contrast, for $m_{1}<15\,{\rm M}_{\odot}$, Pop. II binaries remain the dominant formation channel.
GW signals $\ln\mathcal{D}_{\rm Pop.\,III/Pop.\,II}$
Name $m_{1,\rm s}\,[{\rm M}_{\odot}]$ $m_{2,\rm s}\,[{\rm M}_{\odot}]$ kro1 kro5 lar1 lar5 log1 log2 log3 log4 log5 top1 top5
GW190917 $9.7_{-3.9}^{+3.4}$ $2.1^{+1.1}_{-0.4}$ $-1.1$ $-3.3$ $-0.4$ $-4.2$ $-1.7$ $-3.9$ $-1.3$ $-1.6$ $-4.2$ $-1.1$ $-83.2$
GW191219 $31.1_{-2.8}^{+2.2}$ $1.17_{-0.06}^{+0.07}$ 2.2 3.7 1.9 4.5 2.1 1 2.3 2.2 3.4 1.7 4.1
GW200105 $9.1_{-1.7}^{+1.7}$ $1.91_{-0.24}^{+0.33}$ $-4.4$ $-11.2$ $-4.7$ $-9.9$ $-2.7$ $-16$ $-3.7$ $-3$ $-14.8$ $-3.3$ $-512.4$
Table 7: Detectability-conditioned Bayes factors $\ln\mathcal{D}$ for Pop. III models, evaluated against the fiducial Pop. II model (kro1), for GW events classified as BHNS mergers or containing one component in the lower mass gap. The selected events were detected during the first three observing runs of LVK. Columns 1–3: event name and source-frame component masses. Columns 4–14: $\ln\mathcal{D}$ values for each Pop. III model relative to the Pop. II reference model.
5 Summary
We have investigated the properties of black hole–neutron star binary (BHNS) and binary neutron star (BNS) mergers from Pop. III and Pop. II stars, by means of binary population synthesis simulations performed with sevn (Costa et al. 2023). These merger catalogs encompass the uncertainties on initial mass functions (IMF) and orbital parameters of Pop. III stars.
We have estimated the merger rate densities (MRDs) of both BHNSs (Fig.3) and BNSs (Fig.8) from Pop. III stars and compared them to the MRDs of Pop. II stars.
We find that the MRD of BHNSs from Pop. III stars ranges from $\sim 10^{-2}$ to $2\,\rm Gpc^{-3}\,yr^{-1}$ at redshift $z\sim{13}$. Models with initial orbital parameters from Stacy & Bromm (2013) produce smaller MRDs with respect to models with distributions from Sana et al. (2012). This behavior was already pointed out in Santoliquido et al. (2023) in the case of binary black hole (BH) mergers. The MRD of BHNSs from Pop. II stars dominates over the one from Pop. III stars at least up to $z\sim 17$.
The delay times $t_{\rm del}$ of BHNS mergers for our fiducial models (Fig. 4) follow quite closely the $\propto t^{-1}$ trend from Dominik et al. (2012).
We compare the primary BH mass $m_{1}$ distribution of BHNS mergers for the fiducial models, and find that BHs from Pop. III stars tend to produce a flat distribution in $m_{1}$ with a peak around $22\,{\rm M}_{\odot}$, while BHs from Pop. II stars present a peak at $m_{1}<15\,{\rm M}_{\odot}$ and a decreasing trend up to $\sim 60\,{\rm M}_{\odot}$ (Fig. 5). These different trends in $m_{1}$ result in different trends in the mass ratio $q=m_{2}/m_{1}$ (Fig. 7). BHNSs from Pop. III stars yield a distribution peaking at $q<0.1$ and extending up to $q\sim 0.3$, whereas Pop. II BHNS mergers peak at $q\sim 0.2$ and extend up to $q\sim 0.35$.
Pop. III high-mass BHNSs merge preferentially with large delay times $t_{\rm del}>10^{3}\,\rm Myr$ (Fig. 6). Consequently, second-generation interferometers might be able to detect BHNS mergers with $m_{1}>20\,{\rm M}_{\odot}$ from a metal-free population in the local Universe.
BNSs from Pop. III and II stars mainly merge with short delay times ($t_{\rm del}<50\,\rm Myr$) because of their formation channels, which usually involve one or more common envelope events (Fig. 9). As a result, the peaks of the MRDs of BNSs from both Pop. III and Pop. II stars are close to the peaks of their respective star formation rate density models (Hartwig et al. 2022, 2024; Liu et al., in prep.). We find that the MRDs of metal-free BNSs range from $\sim 10^{-3}$ to $\sim 15\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ at $z\sim 15$. The MRD of Pop. II BNSs peaks at $\mathcal{R}_{\rm BNS}(z)\sim 5\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ around $z\sim 10$, dominating over the Pop. III contribution up to that redshift. Despite the compactness of Pop. III stars during their evolution (Fig. 1), Pop. II stars produce a MRD at least one order of magnitude higher than the Pop. III one. This is mainly due to the impact of the star formation rate density models on the MRDs. Also in this case, configurations with initial orbital parameters from Stacy & Bromm (2013) produce smaller MRDs with respect to those with initial distributions from Sana et al. (2012).
We explore the impact of the initial conditions on the distributions of primary mass $m_{1}$ (Fig. 10) and mass ratio $q$ (Fig. 11) of BHNS mergers from Pop. III stars. We find that all simulated models produce merging BHNSs with $6\lesssim m_{1}\lesssim 50\,{\rm M}_{\odot}$, except for one (log3), which yields BHNS mergers with primary mass up to $400\,{\rm M}_{\odot}$. Models with initial orbital parameters from Stacy & Bromm (2013) create preferentially merging systems with $m_{1}>15\,{\rm M}_{\odot}$, while models with initial orbital parameters from Sana et al. (2012) yield flatter $m_{1}$ distributions with a peak around $22\,{\rm M}_{\odot}$. The impact of the initial orbital configuration is milder for Pop. II stars, because of their largest expansion during their evolution.
We compute the detectability-conditioned Bayes factors (Mould et al. 2023) for three GW signals detected by LVK, classified either as BHNS mergers or as compact mergers with a secondary mass in the lower mass gap ($2-5\,{\rm M}_{\odot}$, Fryer et al. 2012). In our analysis, we compare each simulated Pop. III model to the fiducial Pop. II model (see Table 7). GW191219 stands out as favoring a Pop. III origin, particularly for models that predict mergers with extreme mass ratios. In contrast, GW signals with $m_{1}<20\,{\rm M}_{\odot}$ favor a Pop. II origin.
In summary, our results indicate that, although Pop. III stars are more compact than Pop. II stars, their lower star formation rate density leads to MRDs of BHNSs and BNSs that are at least one order of magnitude lower than for BHNSs and BNSs born from Pop. II binaries. This trend depends on the chosen models for the star formation rate density of Pop. III (Hartwig et al. 2022) and Pop. II stars (Hartwig et al. 2024, Liu et al., in prep.). Finally, we find that Pop. III BHNS mergers can involve massive BHs, and provide a possible formation channel for BHNS mergers with BH mass $>20\,{\rm M}_{\odot}$ detected in the local Universe. The implications of our results for detection rates with second- and third-generation interferometers, as well as the effects of rotation and of different supernova models on the component masses of BHNS mergers, will be explored in future works.
Acknowledgements.
MMa, GC, GI, and FS acknowledge financial support from the European Research Council for the ERC Consolidator grant DEMOBLACK, under contract no. 770017. MMa, RSK and BL also acknowledge financial support from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) STRUCTURES. MB and MMa also acknowledge support from the PRIN grant METE under the contract no. 2020KB33TP. FS acknowledges financial support from the AHEAD2020 project (grant agreement n. 871158). MAS acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101025436 (project GRACE-BH) and from the MERAC Foundation. GC acknowledges partial financial support from European Union—Next Generation EU, Mission 4, Component 2, CUP: C93C24004920006, project ‘FIRES’. GI is supported by a fellowship grant from the la Caixa Foundation (ID 100010434), code LCF/BQ/PI24/12040020. GI also acknowledges financial support under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.4, - Call for tender No. 3138 of 18/12/2021 of Italian Ministry of University and Research funded by the European Union – NextGenerationEU. MMo is supported by the LIGO Laboratory through the National Science Foundation awards PHY-1764464 and PHY-2309200. RSK acknowledges financial support from the ERC via Synergy Grant ”ECOGAL” (project ID 855130) and from the German Ministry for Economic Affairs and Climate Action in project ”MAINN” (funding ID 50OO2206). RSK also thanks the 2024/25 Class of Radcliffe Fellows for highly interesting and stimulating discussions. The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grants INST 35/1597-1 FUGG and INST 35/1503-1 FUGG. We used sevn (https://gitlab.com/sevncodes/sevn) to generate our catalogs (Spera et al. 2019; Mapelli et al. 2020; Iorio et al. 2023). We used $\mathcal{R}$ate (https://gitlab.com/Filippo.santoliquido/cosmo_rate_public) to generate our merger rate densities and convolve our catalogs of mergers with the star formation history (Santoliquido et al. 2020, 2021, 2023). We obtained our star formation rate densities from a-sloth (https://gitlab.com/thartwig/asloth). This research made use of NumPy (Harris et al. 2020), SciPy (Virtanen et al. 2020), Pandas (pandas development team 2020). For the plots we used Matplotlib (Hunter 2007).
Appendix A Initial conditions for Pop. III binary stars
In the following, we briefly present the initial distributions of primary mass, mass ratio, orbital period and eccentricity originally described by Costa et al. (2023) and listed in Table 1.
A.1 Initial mass functions
We consider four different IMFs for the primary mass:
  1. (i)
    A Kroupa (2001) IMF $\xi(m_{\mathrm{ZAMS}})\propto m_{\mathrm{ZAMS}}^{-2.3}$, which is commonly adopted for Pop. II stars. With respect to the canonical IMF, here we consider a single slope since $m_{\textrm{min}}\geq 0.5\,{\rm M}_{\odot}$.
  2. (ii)
    A Larson (1998) IMF $\xi(m_{\mathrm{ZAMS}})\propto m_{\mathrm{ZAMS}}^{-2.35}\exp{\left(-m_{\mathrm{cut1}}/m_{\mathrm{ZAMS}}\right)}$, with $m_{\mathrm{cut1}}=20\,{\rm M}_{\odot}$.
  3. (iii)
    A flat log distribution $\xi(m_{\mathrm{ZAMS}})\propto m_{\mathrm{ZAMS}}^{-1}$ (Stacy & Bromm 2013; Hirano et al. 2015; Susa et al. 2014; Wollenberg et al. 2020; Chon et al. 2021; Tanikawa et al. 2021b; Jaura et al. 2022; Prole et al. 2022).
  4. (iv)
    A top heavy distribution (Stacy & Bromm 2013; Jaacks et al. 2019; Liu & Bromm 2020a), $\xi(m_{\mathrm{ZAMS}})\propto m_{\mathrm{ZAMS}}^{-0.17}\exp{\left(-m_{\mathrm{cut2}}^{2}/m_{\mathrm{ZAMS}}^{2}\right)}$, where $m_{\mathrm{cut2}}=20\,M_{\odot}$.
A.2 Mass ratio and secondary mass
We derive the secondary mass of the ZAMS star with three different prescriptions:
  1. (i)
    The $q$ distribution from Sana et al. (2012), $\xi(q)\propto q^{-0.1}$, with $q\in[0.1,1]$, which fits the mass ratios of O- and B-type binaries in the local Universe; here we set $m_{\mathrm{ZAMS,2}}\geq 2.2\,{\rm M}_{\odot}$.
  2. (ii)
    A ”sorted” distribution: we draw the ZAMS mass of the entire stellar population from the chosen IMF. Then, we randomly pair two stars, enforcing $m_{\mathrm{ZAMS,2}}\leq m_{\mathrm{ZAMS,1}}$.
  3. (iii)
    The $q$ distribution from Stacy & Bromm (2013), $\xi(q)\propto q^{-0.55}$, with $q\in[0.1,1]$. This distribution was obtained by fitting Pop. III stars generated in cosmological simulations. Here we assume $m_{\mathrm{ZAMS,2}}\geq 2.2\,{\rm M}_{\odot}$.
A.3 Orbital period
We assume two distributions for the initial orbital period $P$.
  1. (i)
    $\xi(\Pi)\propto\Pi^{-0.55}$, with $\Pi=\log{(P/\mathrm{day})}\in[0.15,5.5]$ from Sana et al. (2012).
  2. (ii)
    $\xi(\Pi)\propto\exp{\left[-\left(\Pi-\mu\right)^{2}/\left(2\sigma^{2}\right)\right]}$, that is a Gaussian distribution with $\mu=5.5$ and $\sigma=0.85$ (Stacy & Bromm 2013).
A.4 Eccentricity
We draw the orbital eccentricity $e$ from two distributions.
  1. (i)
    $\xi(e)\propto e^{-0.42}$ with $e\in[0,1)$ (Sana et al. 2012).
  2. (ii)
    A thermal distribution $\xi(e)\propto e$ with $e\in[0,1)$ (Heggie 1975; Tanikawa et al. 2021b; Hartwig et al. 2016; Kinugawa et al. 2014), which favors highly eccentric systems. As shown by Park et al. (2021, 2023), Pop. III binaries form preferentially with high orbital eccentricity.
Appendix B BHNS and BNS MRD of Pop. II-I stars
In Figs. 3 and 8, we compare the MRDs of BHNSs and BNSs from Pop. III and Pop. II stars with those from Pop. II-I. These MRDs were derived by Iorio et al. (2023) using a fiducial configuration. Below, we summarize the main properties of this model and describe the computation of the MRD. For further details, we refer the reader to Iorio et al. (2023).
[…]
The resulting BHNS and BNS merger catalogs are used as input for $\mathcal{R}$ate. To compute the MRD of these systems, we use Eq. 2 from Sec. 2.7, assuming the star formation rate density $\psi(z)$ from Madau & Fragos (2017):
$\psi(z)=a\,\frac{(1+z)^{b}}{1+[(1+z)/c]^{d}}\,\rm[{\rm M}_{\odot}\,yr^{-1}\,Mpc^{-3}],$ (4)
where $a=0.01\,{\rm M}_{\odot}\rm\,yr^{-1}\,Mpc^{-3}$ (assuming a Kroupa IMF; Kroupa 2001), $b=2.6$, $c=3.2$, and $d=6.2$.
We also modify Eq. 3 as follows:
$\mathcal{F}(z^{\prime},z,Z)=\frac{1}{M_{\ast}(z^{\prime},Z)}\,\frac{{\rm{d}}\mathcal{N}(z^{\prime},z,Z)}{{\rm{d}}t(z)}\,p(z^{\prime},Z),$ (5)
where we assume the metallicity distribution $p(z^{\prime},Z)$ from Madau & Fragos (2017):
$p(z^{\prime},Z)=\frac{1}{\sqrt{2\pi\sigma_{\rm Z}^{2}}}\exp\left\{-\frac{\left[\log(Z(z^{\prime})/\mathrm{Z_{\odot}})-\langle\log Z(z^{\prime})/\mathrm{Z_{\odot}}\rangle\right]^{2}}{2\,\sigma_{\rm Z}^{2}}\right\}.$ (6)
Here, $\langle\log Z(z^{\prime})/\mathrm{Z_{\odot}}\rangle=\log\langle Z(z^{\prime})/\mathrm{Z_{\odot}}\rangle-\ln(10)\,\sigma_{\rm Z}^{2}/2$, and we adopt $\sigma_{\rm Z}=0.2$ (Bouffanais et al. 2021).
Finally, the total initial stellar mass is computed as $M_{*}(z^{\prime},Z)=M_{\rm sim}/f_{\rm IMF}$, where $M_{\rm sim}$ is the total simulated mass and $f_{\rm IMF}=0.285$, to account for the fact that the simulations only include stars with $m_{\rm ZAMS,1}>5\,{\rm M}_{\odot}$ and $m_{\rm ZAMS,2}>2.2\,{\rm M}_{\odot}$, while the Kroupa IMF is defined down to $0.1\,{\rm M}_{\odot}$ (Kroupa 2001).
Appendix C Model selection
Here, we briefly summarize the method presented by Mould et al. (2023). According to Bayes’ theorem, the posterior distribution of the measured properties of a GW event is given by
$p(\@vec{\theta}|\@vec{d},U)=\frac{\mathcal{L}(\@vec{d}|\@vec{\theta})\,\pi(\@vec{\theta}|U)}{\mathcal{Z}(\@vec{d}|U)},$ (7)
where $\mathcal{L}(\@vec{d}|\@vec{\theta})$ is the likelihood that the vector of waveform parameters ${\theta}$ (e.g., masses, spins, redshifts, etc.) produces the observed data ${d}$, and $\pi(\@vec{\theta}|U)$ is the uninformative prior on the parameters. The normalization factor,
[…]
is the marginal likelihood that normalizes the posterior, representing the probability of observing the data given the chosen uninformative model $U$. Both the prior $\pi(\@vec{\theta}|U)$ and the likelihood $\mathcal{L}(\@vec{d}|\@vec{\theta})$ are conditioned on the uninformative model $U$.
If we include an astrophysical model $A$ as an informative prior, the posterior can be expressed as
$p(\@vec{\theta}|\@vec{d},A)=p(\@vec{\theta}|\@vec{d},U)\,\frac{\pi(\@vec{\theta}|A)}{\pi(\@vec{\theta}|U)}\,\frac{\mathcal{Z}(\@vec{d}|U)}{\mathcal{Z}({\@vec{d}|A)}}.$ (9)
Integrating over ${\theta}$ and rearranging, we obtain the Bayes factor, which quantifies the relative likelihood of the two prior models:
[…]
A value of $\mathcal{B}_{A/U}>1$ ($<1$) implies that the astrophysical model $A$ is more (less) likely than the uninformative prior $U$ to have produced the observed data ${d}$. A decisive conclusion is typically reached when $|\ln\mathcal{B}_{A/U}|\gtrsim 5$ (Jeffreys 1939). If we consider two different informative models, $A$ and $B$, the Bayes factor is simply $\mathcal{B}_{A/B}=\mathcal{B}_{A/U}/\mathcal{B}_{B/U}$.
When comparing two Bayesian models, we can also compute the posterior odds:
$\mathcal{O}_{A/B}=\frac{p(A|\@vec{d})}{p(B|\@vec{d})}=\frac{\pi(A)}{\pi(B)}\,\mathcal{B}_{A/B},$ (11)
which reduces to the Bayes factor $\mathcal{B}_{A/B}$ when the priors $\pi(A)$ and $\pi(B)$ are assumed to be equal, i.e., when we cannot estimate the priors.
We can also account for selection effects associated with the detection of GW signals. When a GW signal is detected, it necessarily originates from the subset of the population that is observable. Therefore, we can generalize Eq. 11 by conditioning the evidence for an astrophysical model $A$ on detectability (det). This leads to:
$\mathcal{Z}(\@vec{d}|A,\mathrm{det})=\frac{P(\mathrm{det}|\@vec{d})\,\mathcal{Z}(\@vec{d}|A)}{P(\mathrm{det}|A)},$ (12)
where the numerator captures information about the specific detected event, and the denominator accounts for the overall distribution of sources that the model predicts to be detectable. This detectability-conditioned evidence ensures that model predictions extending into undetectable regions of parameter space do not artificially improve agreement with observed data. Since $P(\mathrm{det}|\@vec{d})=1$, we can write:
$\frac{\mathcal{Z}(\@vec{d}|A,\mathrm{det})}{\mathcal{Z}(\@vec{d}|U)}=\frac{\mathcal{B}_{A/U}}{P(\mathrm{det}|A)}.$ (13)
Finally, we obtain the detection-weighted Bayes factor:
$\mathcal{D}_{A/B}=\frac{\mathcal{Z}(\@vec{d}|A,\mathrm{det})}{\mathcal{Z}(\@vec{d}|B,\mathrm{det})}=\frac{P(\mathrm{det}|B)}{P(\mathrm{det}|A)}\,\mathcal{B}_{A/B}.$ (14)